# Bootstrapping non-unitary CFTs

### v1

@@ Abstract
In this letter, we present an evolutionary algorithm-based approach to bootstrapping the spectrum of general conformal field theories (CFTs). Starting from a trial spectrum, we invert the crossing equations to extract the corresponding operator product expansion (OPE) coefficients. The statistical distribution of these coefficients, obtained by sampling over cross-ratios, provides a quantitative measure of how closely crossing symmetry is satisfied. We then define a reward function based on this statistic and employ a genetic algorithm to search for spectra that maximize the reward. A key advantage of our framework is that it does not rely on unitarity. We illustrate the method using Virasoro blocks with central charge $c<1$, where the optimal solutions align with the known minimal models. More broadly, this approach - solving systems of nonlinear constraints by exploiting the statistical properties of associated linear variables - provides a general strategy for broad class of bootstrap problems.
@@ I Introduction
Since the seminal work of Rattazzi, Rychkov, Tonni, and Vichi Rattazzi:2008pe, which introduced the modern numerical framework for the conformal bootstrap Polyakov:1974gs - casting crossing symmetry and unitarity as a convex optimization problem and thereby obtaining rigorous bounds on operator dimensions - there has been remarkable progress in constraining the operator spectrum and critical exponents of CFTs across various spacetime dimensions. (see RevModPhys.91.015002 for comprehensive review)
The numerical bootstrap program faces two main challenges. First, casting the CFT bootstrap as an optimization problem naturally maps out the boundary of “theory space” in the space of CFT data. This framework yields powerful constraints when notable theories lie along the boundary, or form isolated “islands” once additional assumptions are imposed—as in the celebrated cases of the 3D Ising and O(N) models PhysRevD.86.025022; El-Showk:2014dwa; Kos:2015mba. Nevertheless, it is highly desirable to explore where inside the theory space do physical theories populate. Second, the reformulation of the bootstrap as an optimization problem relies critically on unitarity. Yet nonunitary CFTs are of considerable interest, appearing in contexts such as RG flows near complex fixed points and holography in more general backgrounds.
To find explicit solutions, even approximately, to the bootstrap equations Belavin:1984vu, one must contend with the challenge of exploring a high-dimensional search space - encompassing conformal dimensions, spins, and OPE coefficients - while satisfying an infinite set of constraints, namely, the vanishing of crossing violations for all values of the conformal cross-ratios. Previous attempts, such as Monte Carlo methods Laio:2022ayq and machine - learning approaches Kantor:2021kbx; K_ntor_2022; Kantor:2022epi, have achieved progress primarily when the search is restricted to a window near known theories.
In this letter, we present a novel strategy that turns these two challenges to our advantage. Given a candidate spectrum that solves the crossing equations, one can invert the constraints to determine the corresponding OPE coefficients. For an approximate spectrum, this inversion yields cross-ratio - dependent variations in the extracted OPE coefficients. The resulting distribution of OPE coefficients over different cross-ratios then provides a quantitative measure of proximity to an exact solution. In this way, the dimensionality of the search space is reduced, while the number of crossing constraints becomes a tunable parameter controlling the statistical quality of the test. This allows us to cast the bootstrap problem into a single objective maximization problem, which can be solved using various numerical approaches.
As a proof of concept, we employ covariant matrix adaptation evolution strategy (CMA-ES) hansen2016cma. For an introduction of CMA-ES as well as the spec for our searches see Appendix C. We illustrate the utility of this method in two-dimensional CFTs, where we solve the bootstrap equation for four scalar primaries, with up to 5 internal states. This represents a 7-dimensional search space. We demonstrate that our algorithm reliably recovers the first three state in the spectrum of known theories, both unitarity and non-unitary.
@@ II The objective function
We consider the four-point function of identical scalar primaries $\phi$ of scaling dimension $\Delta_{\phi}$:
@@
with conformal cross-ratios $\frac{x_{12}^{2}x_{34}^{2}}{x_{13}^{2}x_{24}^{2}}{=}z\bar{z}$, $\frac{x_{14}^{2}x_{23}^{2}}{x_{13}^{2}x_{24}^{2}}=(1{-}z)(1{-}\bar{z})$. The reduced correlator admits the $s$-channel conformal block expansion BELAVIN1984333; Dolan:2000ut
@@
where $f_{\phi\phi\mathcal{O}}$ are the OPE coefficients and $g_{\Delta,s}^{\Delta_{\phi}}(z,\bar{z})$ are the (scalar-external) conformal blocks. The equivalence of $s$- and $t$-channel OPE expansion leads to the crossing equation:
$\sum_{k=1}\big(f_{\phi\phi k}\big)^{2}\,G_{\Delta_{k},s_{k}}^{\Delta_{\phi}}(z,\bar{z})$ $=vac(z,\bar{z}),$ (3)
where $G_{\Delta_{k},s_{k}}^{\Delta_{\phi}}$ is the difference of the $s$ and $t$ channel conformal blocks and $vac(z,\bar{z})$ is the vacuum contribution.
As we will be searching in finite dimensional space, we will be solving the crossing equation with only finite set of spectrum. Therefore it is more fitting to write eq.(3) as:
$\sum_{k=1}^{N}\big(f_{\phi\phi k}\big)^{2}\,G_{\Delta_{k},s_{k}}^{\Delta_{\phi}}(z,\bar{z})+e(z,\bar{z})$ $=vac(z,\bar{z}),$ (4)
where $N$ is the number of low lying states, $e(z,\bar{z})$ represents the contributions from primaries with $\Delta_{i}>\Delta_{N}$. [^1] In practice, there will also be errors coming from the approximation for Virasoro blocks. In the vicinity of the symmetric point $z{=}\bar{z}{=}\frac{1}{2}$, the function $G_{\Delta_{k},s_{k}}^{\Delta_{\phi}}$ is exponentially suppressed for large conformal dimensions Pappadopulo:2012jk. Thus, we expect that for a large class of systems, if we consider points around the symmetric point, $e$ in eq.(4) can be approximately set to zero.
With this in mind, we consider $N$ points $\{z_{j},\bar{z_{j}}\}$ in the vicinity of $z{=}\bar{z}{=}\frac{1}{2}$. For these points eq.(4) can be organized as:
$\sum_{k=1}^{N}C_{k}\,G_{k,j}+e_{j}$ $=v_{j},$ (5)
where $C_{k}=\big(f_{\phi\phi k}\big)^{2}$, $G_{k,j}=G_{\Delta_{k},s_{k}}^{\Delta_{\phi}}(z_{j},\bar{z}_{j})$ and $v_{j}=vac(z_{j},\bar{z}_{j})$. For every choice of the NN $\{z_{j},\bar{z}_{j}\}$ points (we denote them as $\{z\}$) we have a matrix equation, linear in the squared OPE coefficients, $\mathbf{C}$:
@@
Solving the crossing equations with $\mathbf{e}_{\{z\}}$ set to zero leads to OPE coefficients that are formally $z$-dependent, $\mathbf{C}_{\{z\}}$,
@@
If this were the exact bootstrap equation, then the solution $\mathbf{C}_{\{z\}}$ would be in fact $\{z\}$-independent. However, since we are working with an approximate crossing equation, $\mathbf{C}_{\{z\}}$ will depend on $\{z\}$ and their statistical distribution becomes a measure of how close one is to an exact solution. In particular, we can consider $n(\geq N)$ sets of $\{z\}$, each time inverting eq.(6), to obtain $n$ sets of $\mathbf{C}_{\{z\}}$. By taking the standard deviation of the result, we obtain a qualitative measurement on how different $z$ points agree on the values of the OPE coefficients. This standard deviation of the distribution over $\{z\}$ points, which we denote $\sigma_{z}$, is related to the error term, $\mathbf{e}_{\{z\}}$, which we have ignored, since
$\mathbf{C}_{\{z\}}=G_{\{z\}}^{-1}\cdot\mathbf{v}_{\{z\}}=\mathbf{C}+G_{\{z\}}^{-1}\cdot\mathbf{e}_{\{z\}},$ (8)
$\sigma_{z}(\mathbf{C}_{\{z\}})=\sigma_{z}(\mathbf{C}+G_{\{z\}}^{-1}\cdot\mathbf{e}_{\{z\}})=\sigma_{z}(G_{\{z\}}^{-1}\cdot\mathbf{e}_{\{z\}}),$
where $\mathbf{C}$ is the OPE coefficient of the truncated spectrum and independent of $\{z\}$}. We see that the standard deviation of the OPE solutions of the truncated crossing equation, $\sigma_{z}(\mathbf{C}_{\{z\}})$, measures the contribution of the error term.
In practice since $\{z\}$ is distributed in a small neighborhood, the $N\times N$ matrix $G_{\{z\}}$ can become singular and inversion is numerically unstable. To this end we instead sample over $N_{z}\gg N$ number of points and solve for $\mathbf{C}_{\{z\}}$ as the least squares solution. We then consider the following objective function:
$R=-\sum_{i=1}^{N_{r}}\log\left(\left|\frac{\sigma(C_{i\{z\}})}{\text{Mean}(C_{i\{z\}})}\right|\right)\,$, (9)
where $\sigma(C_{i\{z\}})$ is the standard deviation of the solution of the OPE coefficient for the $i$-th state and $N_{r}$ denotes the number of low-lying states whose inferred OPE coefficients are included in the objective. However, when we compare searches with different assumed numbers of internal operators, we will mostly use $N_{r}=1$ and focus on the lowest-dimension state. In that case the reward is directly comparable across spectra with different numbers of internal operators.
For theories where the truncation is a good approximation the reward will be large. As we sample over $N_{z}$ points, the typical relative error (due to statistical fluctuations) of $\sigma(C_{i\{z\}})$ and $\text{Mean}(C_{i\{z\}})$ is $\frac{1}{\sqrt{N_{z}}}$ statistics. Therefore, we expect that this objective function has an associated noise of size $O\left(1/\sqrt{N_{z}}\right)$.
This allows us to instead of scanning through $\{\Delta_{k},s_{k},\big(f_{\phi\phi k}\big)^{2}\}$, as done in for example K_ntor_2022, we only search in the space of $\{\Delta_{k},s_{k}\}$.
@@ III The 2D problem
While our setup is general, in this work we will test it on the scenario: 2d CFTs with only scalar operators. As benchmark we use the A-series minimal models, briefly reviewed in Appendix A. In such case the parameter space include the central charge $c$, conformal dimension of the external $h_{ext}$ and internal operators $h_{int}$. We begin with one internal operator, which is a three-dimensional search, and gradually increase the number of internal operators. For numerical evaluation of the Virasoro blocks Polyakov:1974gs, we utilize the Zamolodchikov-type recursion relations with finite iterations. [^2] We have implemented the recursive algorithm for calculating Virasoro blocks numerically of Chen_2017 on GPU. In this work we used 40 recursion iterations.
We perform the search in four different parameter spaces, characterized by the number of internal states that the known theories in the region contain. In the first three regions, the known theories contain one, two and three internal states respectively. The fourth is characterized by theories with five internal states.
In this work we search by iteratively increase the number of internal state. If at a given number $N$ the found solution falls below a certain threshold, we then add an additional state and restart the search. On the other hand if we have passed the threshold, we still add another state and see if the reward is improved. If it does not and the search finds the same parameters of the first $N$ states with the mean OPE value of the next state close to zero (or more precisely within a standard deviation to zero) we can conclude that $N$ is sufficient. We estimate this threshold by looking at $N_{r}=1$ reward in Eq.(9) for different minimal models. From the result in Fig. 1, we see that it is reasonable to set the threshold of the $N_{r}=1$ reward at 10.
@@
Figure 1: The reward distributions of minimal model processes inside the four regions where we perform the search.
III.1 Search for exact spectra with $c<1$ CFTs
In this part, we perform a search of CFT spectra in regions where the relevant four-point functions admit exact solutions with up to three exchanged Virasoro primaries. For these correlators we include all operators that appear in the OPE channel under consideration, so there is no truncation of the spectrum in Eq. (6). In this case the error term $\mathbf{e}_{\{z\}}$ arises solely from the numerical evaluation of Virasoro blocks, due to the finite number of recursion steps in our implementation.
Two results are presented here. The first, shown in Fig. 2, is global search of theories with a single internal operator exchange. We set the search window to be $c\in\left[-6,1\right]$, $h_{ext}\in\left[-0.4,0.6\right]$ and $h_{int}\in\left[-0.5,1.2\right]$. The full search results and analytic values for minimal models align well as shown in Fig. 2.
@@ III.1 Search for exact spectra with $c<1$ CFTs
Figure 2: Search result in $c<1$ region with single internal operator. The blue dots are search results, while the black ones are analytic values from minimal models.
The second one is a local search in the region where only non-unitary CFTs exist. In particular, we search in the region where $\left(h_{int,1},c,h_{ext}\right)$ are close to that of Yang-Lee CFT. The search windows are given in Table 1 and 2. Note that the relative narrow window of search is due to the large number of minimal models and we aim to test if these models can be identified within this window. We tried to search for theories with $2$ or $3$ internal operators with $h_{int,1}\sim h_{YL}=-0.2$, while keeping the ascending order of $\mathbf{h}_{int}$ ($h_{int,1}\leq h_{int,2}\leq h_{int,3}$). The results are shown in Fig. 3, which again aligns with theoretical values well. In Fig.3, we only displayed the projections onto the plane $\left(h_{int,1},c,h_{ext}\right)$. For those related to $h_{int,2}$ and $h_{int,3}$ are given in Appendix E. Note that in principle at a given $N$, we should find theories with less operator exchange as well. These appear as special cases with part of the OPEs being close to 0. To prevent inefficiency for the search, we put additional penalty for small OPEs.
@@
Table 1: Search window for 22-state search

### v2

@@ Abstract
We introduce a non-unitary-compatible numerical bootstrap strategy based on the statistical stability of OPE data inferred from crossing at multiple cross-ratios. For a trial spectrum, crossing determines OPE coefficients whose residual cross-ratio dependence directly measures the truncation error. This defines a scalar objective on the space of spectra, allowing bootstrap searches without imposing unitarity. Applied to two-dimensional Virasoro blocks, the method reproduces known A-series minimal models, including non-unitary examples, and yields candidate truncated solutions for c>1 with crossing violation comparable to that of minimal models. More generally, our framework provides a practical route to solving bootstrap constraints beyond the convex, unitary setting.
@@ I Introduction
Since the seminal work of Rattazzi, Rychkov, Tonni, and Vichi Rattazzi et al. (2008), the numerical conformal bootstrap has developed into a powerful framework for constraining conformal field theories (CFTs) across spacetime dimensions Polyakov (1974); Poland et al. (2019). By combining crossing symmetry with unitarity, one recasts the bootstrap as a convex optimization problem and obtains rigorous bounds on operator dimensions and OPE coefficients. This perspective has led to remarkably precise results, especially when interesting theories lie on the boundary of the allowed region in CFT data space or form isolated islands under additional assumptions, as in the celebrated examples of the 3D Ising and $O(N)$ models El-Showk et al. (2012, 2014); Kos et al. (2015).
Two complementary challenges nevertheless remain. First, the optimization-based formulation is best suited to delineating the boundary of theory space, whereas it is also desirable to locate theories in its interior. Second, the convex formulation relies crucially on unitarity, while non-unitary CFTs are of broad interest, appearing for example in renormalization-group flows near complex fixed points and in more general holographic settings. This motivates a direct search for solutions to crossing symmetry that does not assume positivity.
This perspective is not new. Gliozzi showed that, given sufficient information about the fusion algebra of low-lying operators, one can detect truncated approximate solutions by analyzing the rank of crossing constraints expanded around the symmetric point Gliozzi (2013); Belavin et al. (1984a). More recently, non-convex search strategies have been explored using Monte Carlo methods Laio et al. (2022) and machine-learning-inspired approaches Kántor et al. (2022b, a, 2023). In practice, however, searching directly over scaling dimensions, spins, and OPE coefficients remains challenging, and successful searches typically require substantial prior input or restriction to neighborhoods of known theories.
In this Letter, we propose a different strategy. For a candidate spectrum, the crossing equations can be inverted to infer the corresponding OPE coefficients. For an exact solution, the inferred OPE data are independent of the choice of cross-ratios. For an approximate or truncated spectrum, by contrast, the inferred OPE coefficients fluctuate as the cross-ratios are varied. The statistical spread of these inferred OPE coefficients therefore provides a direct quantitative measure of proximity to a true solution. [^1] A related use of the spread of inferred OPE data to exclude unphysical spectra appeared in Picco et al. (2016). This observation effectively removes the OPE coefficients from the search space and turns the bootstrap problem into the maximization of a single objective function over the spectrum itself.
As a proof of concept, we apply this idea to two-dimensional CFTs with scalar primaries, using Virasoro conformal blocks and the covariance matrix adaptation evolution strategy (CMA-ES) Hansen (2016) as the optimizer; implementation details are given in Appendix B. The method reproduces known A-series minimal models, including non-unitary examples, in both exact and truncated settings. It also yields stable truncated candidate spectra for non-unitary theories with $c>1$ whose crossing violation is comparable to that of minimal models. These results suggest that the statistical stability of inferred OPE data provides a practical route to bootstrap searches beyond the convex, unitary regime.
@@ II The objective function
We consider the four-point function of identical scalar primaries $\phi$ of scaling dimension $\Delta_{\phi}$,
@@
with cross-ratios defined by $\frac{x_{12}^{2}x_{34}^{2}}{x_{13}^{2}x_{24}^{2}}=z\bar{z}$ and $\frac{x_{14}^{2}x_{23}^{2}}{x_{13}^{2}x_{24}^{2}}=(1-z)(1-\bar{z})$. The reduced correlator admits the conformal block expansion Belavin et al. (1984b); Dolan and Osborn (2001)
@@
where $f_{\phi\phi\mathcal{O}}$ are OPE coefficients and $g_{\Delta,s}^{\Delta_{\phi}}(z,\bar{z})$ are scalar-external conformal blocks. Crossing symmetry may then be written as
$\sum_{k}\big(f_{\phi\phi k}\big)^{2}\,G_{\Delta_{k},s_{k}}^{\Delta_{\phi}}(z,\bar{z})=\mathrm{vac}(z,\bar{z}),$ (3)
where $G_{\Delta_{k},s_{k}}^{\Delta_{\phi}}$ denotes the difference between the $s$- and $t$-channel blocks, and $\mathrm{vac}(z,\bar{z})$ is the vacuum contribution.
In practice we search over truncated spectra with $N$ exchanged operators, so it is convenient to rewrite crossing as
$\sum_{k=1}^{N}\big(f_{\phi\phi k}\big)^{2}\,G_{\Delta_{k},s_{k}}^{\Delta_{\phi}}(z,\bar{z})+e(z,\bar{z})=\mathrm{vac}(z,\bar{z}),$ (4)
where $e(z,\bar{z})$ collects the contribution of omitted operators with $\Delta>\Delta_{N}$. [^2] In practice there is also numerical error from the evaluation of Virasoro blocks. Near the symmetric point $z=\bar{z}={\textstyle\frac{1}{2}}$, the contribution of high-dimension operators is exponentially suppressed Pappadopulo et al. (2012), so for sufficiently well-truncatable spectra the error term can be made small.
We now evaluate Eq. (4) on a set of points $\{(z_{j},\bar{z}_{j})\}$ in a neighborhood of the symmetric point. Writing $C_{k}\equiv\big(f_{\phi\phi k}\big)^{2}$, $G_{k,j}\equiv G_{\Delta_{k},s_{k}}^{\Delta_{\phi}}(z_{j},\bar{z}_{j})$, and $v_{j}\equiv\mathrm{vac}(z_{j},\bar{z}_{j})$, the crossing equations take the matrix form
@@
If we neglect $\mathbf{e}_{\{z\}}$, then for each choice of sampled cross-ratios we infer
@@
For an exact solution, $\mathbf{C}_{\{z\}}$ is independent of the chosen sample. For a truncated or approximate spectrum, however, the inferred OPE coefficients fluctuate with $\{z\}$. Their statistical spread therefore directly measures the residual crossing error:
$\mathbf{C}_{\{z\}}=\mathbf{C}+G_{\{z\}}^{-1}\cdot\mathbf{e}_{\{z\}},\qquad\sigma_{z}(\mathbf{C}_{\{z\}})=\sigma_{z}\!\left(G_{\{z\}}^{-1}\cdot\mathbf{e}_{\{z\}}\right).$ (7)
In practice, the proximity of the sampled points can make $G_{\{z\}}$ ill-conditioned. We therefore take $N_{z}\gg N$ samples and infer $\mathbf{C}_{\{z\}}$ by least squares, which leads to the objective function
$R=-\sum_{i=1}^{N_{r}}\log\!\left|\frac{\sigma(C_{i\{z\}})}{\mathrm{Mean}(C_{i\{z\}})}\right|,$ (8)
where $N_{r}$ counts the low-lying inferred OPE coefficients included in the reward. Large $R$ therefore signals a spectrum whose inferred OPE data are stable under variations of the sampled cross-ratios, and hence approximately satisfy crossing. When comparing truncations with different numbers of exchanged operators, we usually take $N_{r}=1$ and focus on the lowest-dimension exchanged state, making the reward directly comparable across truncations.
The key advantage is that the OPE coefficients are removed from the nonlinear search space: instead of scanning over $\{\Delta_{k},s_{k},(f_{\phi\phi k})^{2}\}$, we search only over $\{\Delta_{k},s_{k}\}$.
@@ III The 2D problem
We test the construction in two-dimensional CFTs with scalar primaries, using Virasoro conformal blocks. As benchmarks we use A-series minimal models, reviewed briefly in Appendix C. The search variables are the central charge $c$, the external conformal weight $h_{\rm ext}$, and the internal weights $h_{{\rm int},i}$. We begin with a single exchanged primary and enlarge the truncation as needed. Virasoro blocks are evaluated numerically using Zamolodchikov-type recursion relations truncated at finite order. [^3] We implemented the recursive algorithm of Ref. Chen et al. (2017) on GPU and use $40$ recursion steps in the present work.
We study four search regions, organized by the number of internal states appearing in the minimal models they contain: one, two, three, and five. The search is iterative. If the best reward for a search with $N$ exchanged states stays below a reference threshold, we repeat the search with $N+1$ states; if the threshold is exceeded, we still test larger truncations and declare the spectrum stabilized only when the reward no longer improves and the newly inferred OPE coefficients are statistically consistent with zero.
To calibrate the threshold, we evaluate the $N_{r}=1$ reward for known minimal models in the four search regions. Figure 1 suggests the empirical benchmark
$R_{N_{r}=1}\simeq 10$ (9)
provides a useful empirical criterion for identifying good approximate solutions.
@@
Figure 1: Reward distributions for minimal-model processes in the four search regions. Their $N_{r}=1$ values motivate the working threshold used below.
III.1 Exact spectra for $c<1$
We first consider regions in which the known solutions involve at most three exchanged Virasoro primaries, so that no physical truncation error is expected. The only residual error then comes from the finite-order evaluation of the Virasoro blocks.
The first search window,
$c\in[-6,1],\qquad h_{\rm ext}\in[-0.4,0.6],\qquad h_{\rm int}\in[-0.5,1.2],$ (10)
contains minimal models with a single exchanged primary. The results, shown in Fig. 2, align closely with the analytic minimal-model values.
@@ III.1 Exact spectra for $c<1$
Figure 2: Single-state search in the $c<1$ region. Blue points are search results and black points are analytic minimal-model values.
We next focus on the region surrounding the Yang-Lee theory. Using narrow search windows centered on that region (listed in Appendix A), the resulting two- and three-state scans, shown in Fig. 3, again agree well with the expected minimal-model spectra; additional projections are collected in Appendix D. These examples show that when the exact spectrum lies within the truncation, the reward function reliably identifies the correct CFT data.
@@
III.2 Truncated spectra for $c<1$
We now turn to a region in which the relevant minimal models contain five exchanged primaries. Starting from a one-state ansatz, we increase the truncation until the best $N_{r}=1$ reward exceeds the reference threshold. The search window and the full reward table are given in Appendix A.
For $N<4$, the reward remains below threshold. At $N=4$, the threshold is exceeded, and the resulting spectra already approximate the first four operators of known five-state processes; representative examples are given in Appendix E. At $N=5$, the reward improves substantially. Increasing the truncation to $N=6$ yields no meaningful improvement, while the inferred OPE coefficient of the sixth state is statistically consistent with zero. Since the first five parameters of the best five- and six-state solutions agree at the percent level, we conclude that the search has stabilized at five states.
To quantify agreement with the exact minimal-model spectra in this seven-dimensional search space, we compare each found point with the nearest minimal-model spectrum using the procedure described in Appendix F. The resulting relative-error distributions are shown in Fig. 4. Relative to the initial random population, CMA-ES sharply improves the agreement with the minimal-model loci, providing a nontrivial validation of the reward function in a genuinely truncated setting.
Figure 4: Relative errors for searches in the region containing five-state processes. CMA-ES substantially sharpens the agreement with nearby minimal-model spectra.
IV Candidate non-unitary spectra with $c>1$
We next apply the method to non-unitary spectra with $c>1$. As before, we begin with a small number of exchanged states and increase the truncation until the $N_{r}=1$ reward exceeds the reference value established from the minimal-model benchmarks.
The initial scan is performed over the broad search window listed in Appendix A. For a $N$-state ansatz, we use the first $N$ entries of the listed ranges for the internal weights together with the corresponding ranges of $c$ and $h_{\rm ext}$. The resulting rewards are shown in Fig. 5: spectra with $N_{r}=1$ reward above threshold first appear at $N=11$.
Figure 5: Initial scan over truncation level in the non-unitary $c>1$ search. Spectra with $N_{r}=1$ reward above threshold first appear at $11$ states.
To test stability under larger truncations, we select a candidate from the $11$-state scan and perform a refined search in a narrower window, again detailed in Appendix A, now allowing for $12$ and $13$ exchanged states. The average $N_{r}=1$ reward increases from $9.86$ at $N=11$ to $10.55$ at $N=12$ and then remains essentially unchanged, $10.53$, at $N=13$, suggesting that the low-lying spectrum has stabilized.
This conclusion is supported by the inferred OPE data. For the highest-reward $13$-state solution (see Table 7 in Appendix A), the inferred OPE coefficients of the $12$th and $13$th states are about three orders of magnitude smaller than that of the $11$th state. Moreover,
$\left|\frac{\sigma(C_{12})}{\mathrm{Mean}(C_{12})}\right|,\qquad\left|\frac{\sigma(C_{13})}{\mathrm{Mean}(C_{13})}\right|,$ (11)
are much larger than the corresponding quantity for the $11$th state, indicating that the extra coefficients are not statistically significant. Since heavier conformal blocks are exponentially suppressed near the symmetric point Pappadopulo et al. (2012), their contribution to the sampled crossing equations is negligible.
A complementary diagnostic is the crossing violation itself. Figure 6 compares the candidate $11$- and $13$-state spectra with minimal models containing $1$, 22, 33, and $5$ exchanged states. Across a broad range of $z$, the candidate $c>1$ spectra have crossing violation comparable to the minimal-model benchmarks, and the similarity of the $11$- and $13$-state curves again indicates that the truncation has stabilized by $11$ states.
Taken together, these observations provide evidence for stable truncated spectra consistent with non-unitary solutions of crossing symmetry at $c>1$. Additional examples are provided in 1, and the five highest-reward $13$-state solutions are listed in Table 8 of Appendix D.
Figure 6: Crossing violation at $\bar{z}=1/2$ for minimal-model benchmarks and for the candidate $11$- and $13$-state $c>1$ spectra. Solid lines denote means and error bars denote standard deviations across spectra with the same truncation.
V Conclusion and outlook
In this work we propose a simple objective function for searching CFT spectra satisfying the crossing equation, that is based on the statistics of solutions of OPE coefficients across many different $z,\bar{z}$ points. This allows us to study solutions to the bootstrap equations for non-unitary CFTs, where conventional convex optimization approaches are no longer applicable. We implemented this framework in two-dimensional CFTs with Virasoro blocks using a fully GPU-parallelized search. In regions where the exact spectrum lies within the truncation, the method reproduces the spectra of A-series minimal models, including non-unitary examples. In regions where truncation is unavoidable, it identifies stable approximate spectra and correctly determines when additional states cease to improve the solution. Extending the search to $c>1$, we found stable truncated candidate spectra whose crossing violation is numerically comparable to that of minimal models.
Several directions deserve further study. Extension to include spin quantum numbers should be straight forward. It will be important to develop a reliable procedure to further test consistency of the spectrum and reconstruct the low lying spectrum. This include extending to mixed correlators and the modularity of torus partition functions Benjamin et al. (2026).
It would also be interesting to improve the treatment of the truncation error $e(z,\bar{z})$ by incorporating analytic expectations for the asymptotic spectrum. In dimensions $d>2$, for example, the light-cone bootstrap implies the existence of large-spin double-twist operators Simmons-Duffin (2017); Kusuki (2019), suggesting a natural way to constrain the tail of the spectrum that is otherwise absorbed into $e$. Since the resulting objective function is structured but generally non-differentiable, it may also provide a useful arena for optimization strategies beyond CMA-ES, including learning-based approaches.
Acknowledgments
We would like to thank Po-Chung Chen, Li-Yuan Chiang, Miranda Cheng, Chong-Sun Chu, David Poland and Ning Su for enlightening discussions. We would especially like to thank Sylvain Ribault for bringing to our attention previous work utilizing the statistical spread of those inferred OPEs. Y-t H thanks Riken iTHEMS and the Yukawa Institute for Theoretical Physics at Kyoto University. Discussions during “Progress of Theoretical Bootstrap” were useful in completing this work. Y-tH is supported by the Taiwan National Science and Technology Council grant 112-2628-M-002-003-MY3 and 114-2923-M-002-011-MY5. S-C Lee is supported by 112-2628-M-002-003-MY3. H Liao is supported in part by the Ministry of Science and Technology grant 112-2112-M-002-024-MY3 and 112-2628-M-002-003-MY3.
Appendix A Supplemental tables
For convenience, we collect here the search windows and representative spectra referred to in the main text.
Table 1: Search window for the $2$-state Yang-Lee-region scan.